# One-Layer Transformers are Provably Optimal for In-context Reasoning and Distributional Association Learning in Next-Token Prediction Tasks

### v2

@@ 5 Empirical Validation
Methods
Full input
to $\xi_{F}$?
Noiseless Noisy
@@ Methods
Origin-Softmax-F
Origin-Softmax-FA
Origin-Linear-F
Origin-Linear-FA
Origin-ReLU-F
Origin-ReLU-FA
Reparam-Softmax-F
Reparam-Softmax-FA
Reparam-Linear-F
[…]
Reparam-Linear-W-FA
Reparam-ReLU-F
[…]
@@
Table 1: Performance comparison of original (no reparameterization) and parameterized models when being trained on the population loss. 0-Loss and Bayes-loss indicate whether a model’s population loss converges to 0 or the Bayes risk on noiseless and noisy settings, respectively. Unseen $y$ indicate whether the performance of a model generalizes to unseen output words. Highlighted methods are both Bayes-optimal and generalize to unseen output words.
@@
Figure 1: Models’ performance when being trained on the population loss for $2000$ iterations. First row, left to right: the population loss, unseen output words test loss and the probability of predicting the output tokens in noiseless learning. Second row, left to right: population loss, unseen output words test loss and the probability of predicting the the noise token in noisy learning with $\alpha=0.8$.
Figure 2: Three models’ population loss and unseen output words test loss when being trained for $100$ epochs on a dataset of $M=2048$ i.i.d sentences, drawn from the noisy data model with $\alpha\in\{0.2,0.8\}$.
@@
To understand the impacts on empirical performances of the reparameterization presented in Sections 3 and 4, we evaluate the one-layer transformers (1) with different choices of parameterization and attention activation functions on the following data model.  [^1] The code for the experiments is available at https://github.com/ngmq/onelayer-transformer-ICR-DA-NTP
Data Model’s Parameters. We set the vocabulary size $N=60$, the embedding dimension $d=128$ and the context length $H=256$. We use $Q=5$ trigger words and $O=4$ output words. The orthogonal embeddings are the standard basis vectors in $\mathbb{R}^{d}$. Note that the unembedding matrix $\bm{U}$} is fixed to be the transpose of the embedding matrix in all of our experiments. For the noisy setting, we choose $\alpha\in\{0.2,0.5,0.8\}$. Our setup generally matches existing works Bietti et al. (2023); Chen et al. (2025).
@@
Our sentences are generated by picking uniformly at random a position for the bigram $(q,y)$ and a position for the bigram $(q,\tau)$ (if the noisy setting). All other words in a sentence are chosen uniformly at random from the set $N]\setminus({\mathcal{Q}}\cup{\mathcal{O}})$. Whenever the population loss is minimized, we run normalized gradient descent with batch size $512$ over $T=2000$ steps. For the finite-sample analysis, we train the models for $100$ epochs on a fixed training set of $M=2048$ samples. The learning rates are in the range $[0.01,0.5]$. Further details and results can be found in Appendix D.
@@
Baselines. We compare $13$ different models which differ in one or more following aspects: model type (Reparam versus Origin), attention type (linear versus ReLU versus Softmax), input to the feed-forward layer (F versus FA) and whether the joint query-key matrix $\bm{W}$} is trained in full without reparameterization. More specifically, the term Origin refers to a model where the three trainable matrices ${\bm{V}},{\bm{W}}$ and $\bm{F}$} are trained without reparameterization, while Param indicates that they are re-parameterized as in Lemmas 3.1 and 4.1. We use FA to denote that the feed-forward layer takes the attention-weighted sum $\sum_{h=1}^{H}\sigma(x_{H}^{\top}{\bm{W}}{\bm{x}}_{h}){\bm{x}}_{h}$ as an input, i.e., $\xi_{F}={\bm{U}}{\bm{F}}({\bm{x}}_{H}+\sum_{h=1}^{H}\sigma(x_{H}^{\top}{\bm{W}}{\bm{x}}_{h}){\bm{x}}_{h})$, and F to denote the opposite. Note that Origin-Softmax-F corresponds to the one-layer transformer in Chen et al. (2025). Finally, the model whose name contains both Reparam and -W has $\bm{V}$} and $\bm{F}$} re-parameterized but $\bm{W}$} is trained in full (as in Theorem 3.4).
5.1 Population Loss Convergence
We train $13$ models, shown in Table 1, to minimize the population loss of noiseless and noisy settings. In the noiseless setting, four models fail to minimize the population loss to 0. Two of these models are re-parameterized with softmax attention, which is due to the fact that the softmax attention compress the attention scores to the range $[0,1]$ and does not allow the logits of output words $\xi_{A,y}$ to be arbitrarily larger than other words in a sentence. This phenomenon can be further observed in Figure 1(c) (and also in Figure 1(f)), where the probability of predicting output tokens (and noise token) by Reparam-Softmax-FA does not converge to the expected value ($1$ for noisless and $\alpha$ for noisy tasks, respectively). On noisy setting, $6$ out of $13$ models fail to minimize the population loss to the Bayes risk. Five of these models use only $x_{H}$ as input to the feed-forward layer, which further indicates that adding the attention-weighted sum $\sum_{h=1}^{H}\sigma(x_{H}^{\top}{\bm{W}}{\bm{x}}_{h}){\bm{x}}_{h}$ plays an important role in achieving Bayes-optimality when training on the whole distribution of the data model. On the other hand, Figures 1(a) and 1(d) show that for the models that converge to Bayes risk, their convergence rate is indeed linear. This observation strongly supports our theoretical claims in Theorems 3.2 and 4.3.
5.2 Generalization on Unseen Output Words
@@
For each model, we examine whether their performance on seen output words in $\mathcal{O}$} is similar to that on unseen output words not in $\mathcal{O}$}. From Table 1, we consistently observe that the ability to generalize to unseen output words is unique to the fully-reparameterized models. In particular, the one-layer transformer loses generalization ability even if $\bm{V}$} and $\bm{F}$} are reparameterized and only $\bm{W}$} is trained from scratch. Figures 1(b) and 1(e) indicate that the test loss on unseen output words even diverges for original, non-reparameterized models. These results strongly suggest that (I) generalization to unseen output words is an important performance criteria for in-context reasoning learning, and (II) while one-layer transformers have the representational capacity to adapt to unseen output words, the solutions found by gradient descent are not naturally biased towards this adaptivity. Thus, further theoretical and empirical studies on this setting are of significant importance to understand emergent abilities in transformers.
5.3 Finite-Sample Training for Noisy Learning
We study the effectiveness of training the models on a small, finite dataset. Figure 2 report the results for three models Reparam-ReLU-FA, Reparam-ReLU-W-FA and Origin-ReLU-FA on $\alpha\in\{0.2,0.8\}$. Note that all of these models use ReLU attention and full-input to the feed-forward layer. Figures 2(a) and 2(c) indicate that after being trained on a finite dataset, all three models successfully generalize to in-distribution sentences (i.e., the output tokens are the same as in the dataset) for both $\alpha<0.5$ and $\alpha>0.5$. However, Figures 2(b) and 2(d) shows a increasing trend in generalization to unseen output words as the models are more restricted to appropriate reparameterization. That is, when all three matrices are trained without reparameterization, they collectively fail to generalize to unseen output words. If only $\bm{V}$} and $\bm{F}$} are reparameterized then the model displays a better, but non-perfect degree of generalization. The fully-reparameterized model achieves perfect generalization, which empirically confirms our generalization result in Theorem 4.4.
@@ 6 Conclusion
We studied the approximation capabilities of one-layer transformer for the challenging in-context reasoning and distributional association learning next-token prediction task. Via a novel reparameterization regime, we rigorously proved that one-layer transformers are capable of achieving Bayes-optimal performance when being trained either directly on the population loss or on a finite dataset. Moreover, the same reparameterization allows one-layer transformers to generalize to sentences that are never seen during training. At the same time, our empirical results also show that without appropriate reparameterization, running gradient descent alone is unlikely to achieve non-trivial out-of-distribution generalization ability. Future works include an in-depth study on the theoretical guarantees and empirical performance of transformers that can simultaneously achieve Bayes-optimality and out-of-distribution generalization to unseen output words.
Appendix A Missing Proofs in Section 3
A.1 Proof of Lemma 3.1
First, we prove the following lemma on the attention scores.
Lemma A.1.
Under Assumption 2.2 and the reparameterization in Lemma 3.1, for all $h\in[H]$ a sentence with trigger word $q$, we obtain ${\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h}=\lambda_{q}\mathbb{1}\{z_{h-1}=q\}$.
@@ Appendix A Missing Proofs in Section 3
Let ${\bm{W}}_{|k}=\lambda_{k}E(k)\tilde{E}^{\top}(k)$. By Assumption 2.2, we have
@@
${\bm{x}}_{H}^{\top}W{\bm{x}}_{h}$ $=\sum_{k\in{\mathcal{Q}}}{\bm{x}}_{H}^{\top}W_{|k}{\bm{x}}_{h}=\lambda_{q}\mathbb{1}\{z_{h-1}=q\}.$
Lemma A.1 indicates that the attention scores are always non-negative. As a result, for both linear and ReLU attention, we have $\sigma({\bm{x}}_{H}^{\top}W{\bm{x}}_{h})={\bm{x}}_{H}^{\top}W{\bm{x}}_{h}$. Hence, it suffices to prove Lemma 3.1 and our subsequent results for linear attention. More generally, our proof can be extended to any activation function where $\sigma(x)=cx$ for $c>0,x\geq 0$.
@@
(Of Lemma 3.1) Fix a trigger token $q\in{\mathcal{Q}}$ and an output token $y\in{\mathcal{O}}$. Consider sentences that contain $q$ and $y$ as their trigger and output, respectively. By Lemma A.1 and Assumption 2.2, for the linear attention model, we have
@@
[…]
$L(\bm{\lambda}_{t})=O\left(\ln(1+\frac{N-1}{\exp(t)})\right)=O(N\exp(-t)).$
A.3 Proof of Theorem 3.3
@@
Since Lemma A.1 holds for any set of output tokens, replacing $y$ by $y_{\mathrm{test}}$ everywhere does not affect the attention scores ${\bm{x}}_{H}{\bm{W}}{\bm{x}}_{h}=\lambda_{q}\mathbb{1}\{z_{h-1}=q\}$. This implies that by Equation 9, we have $\xi_{j}=\xi_{A,j}=\lambda_{q}C_{q,j}\mathbb{1}\{j=y_{\mathrm{test}}\}$. Therefore,
@@
A.4 Directional convergence of running gradient descent on the joint query-key matrix
@@
otherwise. Taking the differential on both sides, we obtain

### v3

@@ 6 Empirical Validation
Models Noiseless Noisy
@@
Origin-Linear
Origin-ReLU
Origin-Softmax
Reparam-Linear-W
Reparam-ReLU-W
Reparam-Softmax-W
Reparam-Softmax
Reparam-Linear
Reparam-ReLU
@@
Table 1: Performance comparison of original (no reparameterization) and parameterized models trained on the population loss. 0-Loss and Bayes-loss indicate whether a model’s Bayes-optimal in noiseless and noisy settings, respectively. Unseen $y$ indicate whether a model generalizes to unseen output words. The best models are highlighted.
@@
Figure 1: Population and OOD test loss of models trained on the population loss in noiseless and noisy settings. First row, left to right: the population loss, OOD test losses of Origin (no parameterization) models and OOD test losses of re-parameterized models in noiseless learning. Second row, left to right: the corresponding losses as in the first row in the noisy setting with $\alpha=0.5$.
@@
To understand the impacts on empirical performances of the reparameterization presented in Sections 3 and 5, we evaluate the one-layer transformers (1) with different choices of parameterization and attention activation functions on the following data model.
@@
Data Model’s Parameters. We set the vocabulary size $N=60$, the embedding dimension $d=128$ and the context length $H=256$. We use $Q=5$ trigger words and $O=4$ output words. The orthogonal embeddings are the standard basis vectors in $\mathbb{R}^{d}$. For the noisy setting, we choose $\alpha\in\{0.2,0.5,0.8\}$ to cover all three cases of $\alpha<0.5,\alpha=0.5$ and $\alpha>0.5$. Our sentences are generated by picking uniformly at random a position for the bigram $(q,y)$ and a position for the bigram $(q,\tau)$ (in the noisy setting). All other words in a sentence are chosen uniformly at random from the set $N]\setminus({\mathcal{Q}}\cup{\mathcal{O}})$. For training on the population loss, we run normalized gradient descent with batch size $512$ over $T=2000$ steps. For the finite-sample analysis, we train the models for $100$ epochs on a fixed training set of $M=2048$ samples. Further details and results are in Appendix F.
@@
Baselines. We compare $9$ different models which differ in one or more following aspects: model type (Reparam versus Origin), attention type (linear versus ReLU versus Softmax), and whether the joint query-key matrix $\bm{W}$} is trained in full without reparameterization. More specifically, the term Origin refers to a model where the three trainable matrices ${\bm{V}},{\bm{W}}$ and $\bm{F}$} are trained without reparameterization, while Param indicates that they are re-parameterized as in Lemmas 3.14.1 and 5.1. Note that Origin-Softmax corresponds to the one-layer transformer in Chen et al. (2025). Finally, the model whose name contains both Reparam and -W has $\bm{V}$} and $\bm{F}$} re-parameterized but $\bm{W}$} is trained in full without any reparameterization.
6.1 Loss Convergence in Noiseless and Noisy Learning
We train $9$ models, shown in Table 1, to minimize the population loss of noiseless and noisy settings. In the noiseless setting, only Origin-Linear fails to achieve a zero loss, indicating the limited approximation power of linear attention. In noisy setting, $4$ out of $9$ models fail to achieve the Bayes risk. These four models consists of the three Origin models and Reparam-Linear-W. This shows that despite the high expressive power of one-layer transformers, gradient descent alone may not be able to find the in-distribution optimal solution. In contrast, all fully-parameterized one-layer models, including the one with linear attention, converge to Bayes-optimal solutions. This emphasizes the crucial role of structural parameterization in guiding gradient descent training towards better solutions, especially on models with low expressive power (i.e. linear attentions). In addition, Figures 1(a) and 1(d) show that for the models that converge to Bayes risk, their convergence rate is indeed linear. This empirically supports our theoretical claims in Theorems 3.24.2 and 5.4.
6.2 OOD Generalization on Unseen Output Words
@@
For each model, we examine whether their performance on seen output words in $\mathcal{O}$} is similar to that on unseen output words not in $\mathcal{O}$}. Table 1 shows that that the ability to generalize to unseen output words consistently increase with more parameterization and a model’s expressiveness. On the other hand, for models with limited expressive power (i.e. one-layer linear transformers), parameterization plays a more important role. In particular, when all three matrices are trained without reparameterization, they collectively fail to generalize to unseen output words regardless of the type of attention. Figures 1(b) and 1(e) indicate that the test loss on unseen output words even diverges for original, non-reparameterized models. These results strongly suggest that (I) generalization to unseen output words is an important performance criteria for in-context reasoning learning, and (II) while one-layer transformers have the representational capacity to adapt to unseen output words, the solutions found by gradient descent are not naturally biased towards this adaptivity. On the other hand, Figures 1(c) and 1(f) shows that combining partial parameterization (on $\bm{V}$} and $\bm{F}$}) and non-linear attention functions (e.g. ReLU or softmax) already leads to models that are simultaneously Bayes-optimal and generalize out-of-distribution. Moreover, similar to Bayes-optimality, OOD generalization can also be obtained with linear attention by careful parameterization. These empirical results hold for both $\alpha\leq 0.5$ and $\alpha>0.5$, which further confirms the generality of our Theorem 5.5.
@@ 7 Conclusion
We studied the approximation capabilities of transformer for an one-step in-context recall task. Via a novel reparameterization regime, we rigorously proved that one-layer transformers are capable of achieving Bayes-optimal performance when being trained either directly on the population loss or on a finite dataset. Moreover, the same reparameterization allows one-layer transformers to generalize to sentences that are never seen during training. At the same time, our empirical results also show that without appropriate reparameterization, running gradient descent alone is unlikely to achieve non-trivial out-of-distribution generalization ability. Future works include an in-depth study on the theoretical guarantees and empirical performance of non-parameterized transformers that can simultaneously achieve Bayes-optimality and out-of-distribution generalization.
Appendix A Additional example sentences for the in-context recall task
The in-context recall task considered in our paper encompasses a large range of practical linguistic scenarios. In this section, we provide additional example sentences in two domains: object identification and transitive inference. In all of these examples, each sentence contains at least one bigram $(q,y)$ before the last query word.
A.1 Object Identification
The task is to identify the right in-context object. Examples include:
Input: “To Harry” were the first two words in a letter that Ron and Hermione wrote to [?]
Output: Harry.
Input: People living the province of Quebec are proud of the natural beauty of the [?]
Output: province.
Input: You should travel on Sunday instead of on Monday, since there is a lot of traffic on [?]
Output: Monday.
A.2 Transitive Inference
The task is to identify the right object that has a specific relationship with other objects in the sentence. Examples include:
Input: If London is on the same continent as Paris, and Paris is on the same continent as Milan, then London is on the same continent as [?]
Output: Milan.
Input: If the table has the same color as the book, and the book has a different color than the chair, then the table has a different color than the [?]
Output: chair.
Input: If the GDP of Germany is larger than the combined GDP of Singapore and Spain, then it is certain that the GDP of Spain is smaller than the GDP of [?]
Output: Germany.
Appendix B Missing Proofs in Section 3
B.1 Proof of Lemma 3.1
First, we prove the following lemma on the attention scores. We write $\{a=b=c\}$ for the event that $a,b,$ and $c$ are equal, i.e. $\{a=b\cap b=c\}$.
Lemma B.1.
Under the reparameterization in Lemma 3.1, for all $h\in[H]$ a sentence with trigger word $q$, we obtain ${\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h}=\lambda_{q}\mathbb{1}\{z_{h-1}=z_{H}=q\}$.
@@ Appendix B Missing Proofs in Section 3
Let ${\bm{W}}_{|k}=\lambda_{k}E(k)\tilde{E}^{\top}(k)$. We have
@@
${\bm{x}}_{H}^{\top}W{\bm{x}}_{h}$ $=\sum_{k\in{\mathcal{Q}}}{\bm{x}}_{H}^{\top}W_{|k}{\bm{x}}_{h}=\lambda_{q}\mathbb{1}\{z_{h-1}=z_{H}=q\}.$
Lemma B.1 indicates that the attention scores are always non-negative. As a result, for both linear and ReLU attention, we have $\sigma({\bm{x}}_{H}^{\top}W{\bm{x}}_{h})={\bm{x}}_{H}^{\top}W{\bm{x}}_{h}$. Hence, it suffices to prove Lemma 3.1 and our subsequent results for linear attention. More generally, our proof can be extended to any activation function where $\sigma(x)=cx$ for $c>0,x\geq 0$.
@@
(Of Lemma 3.1) Fix a trigger token $q\in{\mathcal{Q}}$ and an output token $y\in{\mathcal{O}}$. Consider sentences that contain $q$ and $y$ as their trigger and output, respectively. By Lemma B.1, for the linear attention model, we have
@@
$\frac{\partial L}{\partial\lambda_{1,t}}$ $=\mathbb{E}_{y,z}\left[C_{q_{1},y}\frac{\exp(C_{q_{1},y}\lambda_{1,t})}{\exp(C_{q_{1},y}\lambda_{1,t})+N-1}-1\right]$
$=\mathbb{E}_{y,z}\left[C_{q_{1},y}\frac{\exp(C_{q_{1},y}\lambda_{k,t})}{\exp(C_{q_{1},y}\lambda_{k,t})+N-1}-1\right]$
$=\mathbb{E}_{y,z}\left[C_{q_{k},y}\frac{\exp(C_{q_{k},y}\lambda_{k,t})}{\exp(C_{q_{k},y}\lambda_{k,t})+N-1}-1\right]$
$=\frac{\partial L}{\partial\lambda_{k,t}},$
@@
$L(\bm{\lambda}_{t})=O\left(\ln(1+\frac{N-1}{\exp(t)})\right)=O(N\exp(-\eta t)).$
B.3 Proof of Theorem 3.3
@@
Since Lemma B.1 holds for any set of output tokens, replacing $y$ by $y_{\mathrm{test}}$ everywhere does not affect the attention scores ${\bm{x}}_{H}{\bm{W}}{\bm{x}}_{h}=\lambda_{q}\mathbb{1}\{z_{h-1}=q\}$. This implies that by Equation 5, we have $\xi_{j}=\xi_{A,j}=\lambda_{q}C_{q,j}\mathbb{1}\{j=y_{\mathrm{test}}\}$. Therefore,
@@
B.4 Directional convergence of running gradient descent on the joint query-key matrix
@@
Taking the differential on both sides, we obtain